Consider a solenoid that is 50 cm long and has a cross-sectional area of 5 cm2. It is wound with 400 turns of wire, and a current of 2.3 A is supplied to it. Assume that its core has a relative permeability of 400. How much flux is in the solenoid?


1
Expert's answer
2022-04-11T12:07:06-0400

The magnetic field inside a solenoid

"B=\mu n I=k\mu_0\frac{N}{L}I"

"B=400*4\pi*10^{-7}*\frac{400}{0.5}*2.3=0.924\:\rm T"

The magnetic flux

"\Phi=BA\\
=0.924*5*10^{-4}=4.6*10^{-4}\:\rm Wb"